# Radio Map-Based Spectrum Sharing for Joint Communication and Sensing

### v1

@@ III-A PROBLEM FORMULATION
where (10b) is the ergodic rate requirement of the communication user, (10c) and (10d) are the maximal transmit power constraints of BSs and radars, and (10e) and (10f) are the sum-power constraints of C&S systems.
@@ III-B1 CLOSED-FORM APPROXIMATION
(P1) is highly complex due to the non-convexity of the objective and constraint. The most challenging aspect is the ergodic rate expression (8), which is encapsulated within the expectation operation and lacks a closed-form representation. To address this challenge, we first introduce a determined approximation to $\bar{R}(\mathbf{P}_{c},\mathbf{P}_{r})$ [35]:
@@
The aforementioned approximation is derived under the assumption that transmitters and receivers are equipped with an infinite number of antennas. In this limit, the random channel matrix tends to be deterministic. Leveraging the random matrix theory, we can extract the expectation operation and derive this approximation. Readers can refer to [35] for a detailed proof. On this basis, we transform (P1) into the following problem,
@@ III-B2 AUXILIARY-FUNCTION-BASED SCALING
{$*.\displaystyle\Rightarrow\left\{\begin{split}&g(\mathbf{P}_{c},\mathbf{P}_{r},z)\geqslant R_{req},\\ &1\leqslant z\leqslant v^{*}.\end{split}\right.$
@@
where $\gamma$ is an introduced auxiliary variable. Since $\rho_{i}(\mathbf{P}_{r},\mathbf{P}_{c})$ is a ratio expression of $\mathbf{P}_{c}$ and $\mathbf{P}_{r}$, (30b) can be tackled by the fraction programming scheme named quadratic transform [36],
$\max\limits_{\beta_{i}}\bigg{[}2\beta_{i}\sqrt{[h_{{r_{i}},i}]^{H}p_{r_{i}}h_{{r_{i}},i}}-\beta_{i}^{2}\sigma^{2}_{r_{i}}(\mathbf{P}_{c})\bigg{]}\geqslant\ \gamma,i=1,...,M_{r},$ (31)
@@
$\beta_{i}^{*}=\frac{\sqrt{[h_{r_{i},i}]^{H}p_{r_{i}}h_{r_{i},i}}}{\sigma^{2}_{r_{i}}(\mathbf{P}_{c})},i=1,...,M_{r}.$ (32)
@@
$2(\beta_{i}^{*})^{s-1}\sqrt{[h_{{r_{i}},i}]^{H}p_{r_{i}}h_{{r_{i}},i}}-((\beta_{i}^{*})^{s-1})^{2}\sigma^{2}_{r_{i}}(\mathbf{P}_{c})\geqslant\ \gamma.$ (35)
@@
$2(\beta_{i}^{*})^{s-1}\sqrt{[h_{{r_{i}},i}]^{H}p_{r_{i}}h_{{r_{i}},i}}$
@@
where $\mathcal{G}(z,\mathbf{t}|\mathbf{t}^{s-1})$ and $\mbox{In}\big{(}t_{j}|t_{j}^{s-1}\big{)}+\mbox{In}\big{(}z|z^{s-1})\big{)}+\mbox{In}\big{(}\sigma_{c}^{2}(\mathbf{P}_{r})|\sigma_{c}^{2}((\mathbf{P}_{r})^{s-1})\big{)}$ are the Taylor expansion expressions of $\mathcal{G}(z,\mathbf{t})$ and $\mbox{In}(t_{j})+\mbox{In}(z)+\mbox{In}(\sigma_{c}^{2}(\mathbf{P}_{r}))$. Their expressions are given at the top of the this page. The detailed algorithm is summarized in Algorithm 1. The proof of convergence is given in Appendix B.
@@
IV SIMULATION RESULTS AND DISCUSSIONS
In this section, we present simulation results and discussions. The simulation layout is the same as that we established in the case study of the radio map, as presented in Fig. 4 (a). The layout includes three BSs, i.e., $M_{c}=3$, two radars, i.e., $M_{r}=2$, one interfered user, and a red diamond-shaped target. They are located at (604 m, 629 m), (1289 m, 2022 m), (1986 m, 1316 m), (-1167 m, 3125 m), (2620 m, -779 m), (650 m, 1134 m), and (1360 m, 1000 m), respectively. BSs and the interfered user are equipped with half-wave dipole antennas. The antenna number of the user is equal to the number of BSs, i.e, $N_{c}=3$. Radars are equipped with directional antennas. The maximum antenna gain is $30$ dBi and the half-power beam width, denoted as $\theta_{\text{3dB}}$, is $32^{\circ}$. The carrier frequency is $f_{c}=2.8$ GHz, which is the frequency band of the surveillance radar, and the noise variance is $\sigma^{2}=-107\ \text{dBm}$ [32]. Power related parameters are set as $P_{cmax}=40$ W, $P_{csum}=100$ W, $P_{rmax}=1$ kW and $P_{rsum}=1.5$ kW. This is because radars have much higher power than BSs [24]. The sampling number of radars is $N=512$ per CPI, the false alarm probability is $P_{F}=1\times 10^{-4}$, and thus, $\mu=0.0182$ according to (6).
@@ IV SIMULATION RESULTS AND DISCUSSIONS
We first evaluate the impact of large-scale CSI accuracy on C&S performance. As shown in Fig. 5, the curve corresponding to the radio-map-predicted large-scale CSI closely aligns with the one using the real data, while the curve obtained through the curve-fitting method exhibits a noticeable gap. This discrepancy is expected, given that the radio map provides a more accurate estimation of the real data, as demonstrated in Fig. 4. In our case study, the average deviations of the large-scale CSI map generated by the radio map are below $4$ dB while the deviations associated with the curve-fitting method exceed $25$ dB. Such a substantial deviation leads to misinformed power allocation and, consequently, significant performance deterioration. Therefore, compared with the generic channel model, which is blind to the surrounding environment, the radio map verifies the importance of being environmental-aware. In addition, since the radio map does not require the pilot-based C&S interactions, it is cost-effective for practical applications.
In Fig. 6, we compare the proposed scheme with the scheme proposed in [23] and the unilateral schemes of optimizing $\mathbf{P}_{c}$ or $\mathbf{P}_{r}$ only. Since the scheme proposed in [23] considered the power allocation between one BS and one radar, we used the equal power allocation among different BSs and radars for a fair comparison. The equal power allocation was also used on the not optimized side in the simulation of the unilateral schemes. The results demonstrate that, under the same ergodic rate requirement, the proposed scheme achieves the highest detection probability. Leveraging the large-scale CSI, the proposed scheme intelligently allocates more power to the BSs and radars with favorable serving channels and unfavorable interference channels. This amplifies the positive effects of power for their own systems while constraining the negative effects on the opposite system. As a result, the interference between C&S systems is significantly decoupled in the spatial domain. In contrast, the scheme proposed in [23] is blind to the large-scale CSI, treating all BSs and radars equally in the power allocation. This uniform allocation can not avoid signal collisions, especially when the transmitting equipment on one side and the receiving equipment on the other side are in close proximity. As for unilateral schemes, their efficacy is constrained by the limited DoFs. Only optimizing $\mathbf{P}_{c}$ even failed to provide solutions when $R_{req}\geqslant 6$ bits/s/Hz in the considered setting. The is because when the sensing system does not cooperate to control its power, the interfered user is exposed to a strong-interference environment. Under the given power constraints, the ergodic user rate is limited to below $6$ bits/s/Hz, regardless of how the communication power is allocated. In this sense, the joint power allocation is more effective to provide scalable C&S services.
@@
In Fig. 7, we show the impact of the power budgets, i.e., $P_{csum}$ and $P_{rsum}$, on C&S performance. It is evident that the detection performance improves with the availability of the power budgets. In addition, we find that the curves under different $P_{csum}$ overlap when the ergodic rate requirement of the interfered user, i.e., $R_{req}$, is low, and the curves under different $P_{rsum}$ overlap when $R_{req}$ is high. The reason is that, when $R_{req}$ is low, the communication side can meet the ergodic rate requirement without utilizing its full power. To maximize the radar SINR, the communication side conservatively reserves power and the sensing side operates at full capacity. Consequently, the curves under different $P_{csum}$ overlap. In contrast, when $R_{req}$ is high, the sensing side has to control its power and the communication side works at full capacity. This leads to the curves of different $P_{rsum}$ overlapping in the high $R_{req}$ region. We further find that, in spectrum sharing scenarios where two systems are mutually restricted, the optimal solution lies at the boundary of the feasible set but not within it. In other words, one system must exhaust its resources in the optimal solution.
@@
In Fig. 8, we show the impact of radar antenna directivity, depicted by the half-power beam width, i.e., $\theta_{\text{3dB}}$, on the system performance. It can be seen that the detection probability experiences serve degradation as the half-power beam width increases. This is because the larger beam width leads to more power leakage to the communication user and also opens a larger hole to receive BS interference. In addition, the results show that the gap between the proposed scheme and the scheme proposed in [23] is more obvious when $\theta_{\text{3dB}}$ is large. The reason is that, when $\theta_{\text{3dB}}$ is small, the C&S signals are largely decoupled by directional beams. The improved space left for the proposed scheme is restricted [23]. In contrast, when $\theta_{\text{3dB}}$ is large, C&S signals are mutually coupled. The proposed scheme functions like directive beams to separate C&S signals into a comb-like pattern. Consequently, it significantly outperforms the scheme proposed in [23]. This reveals that the proposed large-scale-CSI-based scheme works in the spatial domain, essentially optimizing the spatial distribution of power to mitigate C&S interference.

### v2

@@ III-A PROBLEM FORMULATION
where (13b) is the ergodic rate requirement of the interfered user, (13c) and (13d) are the maximal transmit power constraints of BSs and radars, and (13e) and (13f) are the sum-power constraints of C&S systems. In addition, while the proposed scheme primarily takes the radar SINR as the objective, the communication performance is preferentially satisfied as we consider it in the constraint. In practical implementations, the ergodic rate threshold is adjusted to meet the requirement of the interfered user. However, this needs manual adjustments of the threshold parameter to provide satisfying C&S solutions for different scenarios. To achieve a more adaptable balance between C&S functions, a compound metric that can uniformly measure C&S performance is necessary. We call for more research attention to this direction, which would present exciting revenues for the further development of JCAS.
@@ III-B1 CLOSED-FORM APPROXIMATION
(P1) is highly complex due to the non-convexity of the objective and constraint. The most challenging aspect is the ergodic rate expression (12), which is encapsulated within the expectation operation and lacks a closed-form representation. To address this challenge, we first introduce an approximation to $\bar{R}(\mathbf{P}_{c},\mathbf{P}_{r})$ [53]:
@@
The aforementioned approximation is derived under the assumption that transmitters and receivers are equipped with an infinite number of antennas. In this limit, the random channel matrix tends to be determined. Leveraging the random matrix theory, we can extract the expectation operation and derive this approximation. Readers can refer to [53] for a detailed proof. On this basis, we transform (P1) into the following problem,
@@ III-B2 AUXILIARY-FUNCTION-BASED SCALING
$\Rightarrow\left\{\begin{split}&g(\mathbf{P}_{c},\mathbf{P}_{r},z)\geqslant R_{req},\\ &1\leqslant z\leqslant v^{*}.\end{split}\right.$
@@
where $\gamma$ is an introduced auxiliary variable. Since $\rho_{i}(\mathbf{P}_{r},\mathbf{P}_{c})$ is a ratio expression of $\mathbf{P}_{c}$ and $\mathbf{P}_{r}$, (33b) can be tackled by the fractional programming scheme named quadratic transform [54],
$\max\limits_{\beta_{i}}\bigg{[}2\beta_{i}\sqrt{[h_{{r_{i}}\rightarrow{r_{i}}}]^{H}p_{r_{i}}h_{{r_{i}}\rightarrow{r_{i}}}}-\beta_{i}^{2}\sigma^{2}_{r_{i}}(\mathbf{P}_{c})\bigg{]}\geqslant\ \gamma,$ (34)
@@
$\beta_{i}^{*}=\frac{\sqrt{[h_{{r_{i}}\rightarrow{r_{i}}}]^{H}p_{r_{i}}h_{{r_{i}}\rightarrow{r_{i}}}}}{\sigma^{2}_{r_{i}}(\mathbf{P}_{c})},i=1,...,M_{r}.$ (35)
@@
$2(\beta_{i}^{*})^{s-1}\sqrt{[h_{{r_{i}}\rightarrow{r_{i}}}]^{H}p_{r_{i}}h_{{r_{i}}\rightarrow{r_{i}}}}-((\beta_{i}^{*})^{s-1})^{2}\sigma^{2}_{r_{i}}(\mathbf{P}_{c})\geqslant\ \gamma.$ (38)
@@
$2(\beta_{i}^{*})^{s-1}\sqrt{[h_{{r_{i}}\rightarrow{r_{i}}}]^{H}p_{r_{i}}h_{{r_{i}}\rightarrow{r_{i}}}}$
@@
where $\mathcal{G}(z,\mathbf{t}|\mathbf{t}^{s-1})$ and $\mbox{In}\big{(}t_{j}|t_{j}^{s-1}\big{)}+\mbox{In}\big{(}z|z^{s-1})\big{)}+\mbox{In}\big{(}\sigma_{c}^{2}(\mathbf{P}_{r})|\sigma_{c}^{2}((\mathbf{P}_{r})^{s-1})\big{)}$ are the Taylor expansion expressions of $\mathcal{G}(z,\mathbf{t})$ and $\mbox{In}(t_{j})+\mbox{In}(z)+\mbox{In}(\sigma_{c}^{2}(\mathbf{P}_{r}))$. Their expressions are given at the top of the this page. The detailed algorithm is summarized in Algorithm 1. The proof of convergence is given in Appendix B. In addition, we analyze the computation complexity of the proposed algorithm. (P6) is a convex problem. If the classical interior point method is employed to solve it, the computational complexity is given by $\mathcal{O}\left(({2M_{c}+M_{r}+2})^{3.5}\log(1/\epsilon)\right)$, where $(2M_{c}+M_{r}+2)$ represents the dimension of the optimization variables and $\epsilon_{0}$ is the target accuracy. Assuming that the proposed algorithm converges within $I$ iterations, the total computational complexity is thus $\mathcal{O}\left(I({2M_{c}+M_{r}+2})^{3.5}\log(1/\epsilon)\right)$.
Figure 5: The number of iterations of the proposed iterative algorithm under 100 randomly generated typologies.
IV SIMULATION RESULTS AND DISCUSSIONS
@@
In this section, we present simulation results and discussions. The simulation layout is the same as that we established in the case study of the radio map, as presented in Fig. 4 (a). The layout includes three BSs, i.e., $M_{c}=3$, two radars, i.e., $M_{r}=2$, one interfered user, and a red diamond-shaped target. They are located at (604 m, 629 m), (1289 m, 2022 m), (1986 m, 1316 m), (-1167 m, 3125 m), (2620 m, -779 m), (650 m, 1134 m), and (1360 m, 1000 m), respectively. BSs and the interfered user are equipped with half-wave dipole antennas. The antenna number of the user is equal to the number of BSs, i.e, $N_{c}=3$. Radars are equipped with directional antennas. The maximum antenna gain is $30$ dBi and the half-power beam width, denoted as $\theta_{\text{3dB}}$, is $32^{\circ}$. The carrier frequency is $f_{c}=2.8$ GHz, which is the frequency band of the surveillance radar, and the noise variance is $\sigma^{2}=-107\ \text{dBm}$ [46]. Power-related parameters are set as $P_{cmax}=40$ W, $P_{csum}=100$ W, $P_{rmax}=1$ kW and $P_{rsum}=1.5$ kW. This is because radars have much higher power than BSs [31]. The sampling number of radars is $N=256$ per CPI, the false alarm probability is $P_{F}=1\times 10^{-4}$, and thus, $\mu=0.0182$ according to (8).
We first present the convergence performance of the proposed iterative algorithm. In this simulation, we randomly changed the user and target positions and ran the proposed algorithm. The number of iterations was recorded and presented in Fig. 5. Remarkably, in over $90\%$ cases, the proposed algorithm converges within 10 times. The maximum number of iterations observed is only $12$. These results affirm the rapid convergence of the proposed algorithm.
Figure 7: Comparisons between the large-scale-CSI-based scheme and the full-CSI-based scheme
@@
We evaluate the impact of large-scale CSI accuracy on C&S performance in Fig. 6. It can be seen that the curve corresponding to the radio-map-predicted large-scale CSI closely aligns with the one using the real data, while the curve obtained through the curve-fitting method exhibits a noticeable gap. This discrepancy is expected, given that the radio map provides a more accurate estimation of the real large-scale CSI, as demonstrated in Fig. 4. In our case study, the average deviation of the large-scale CSI map generated by the radio map is below $4$ dB while the deviation associated with the curve-fitting method exceeds $25$ dB. Such a substantial deviation leads to misinformed power allocation and, consequently, performance deterioration. Therefore, compared with the channel model, which is blind to the surrounding environment, the radio map verifies the extrinsic information, i.e., positions and surroundings, is effective in coordinating conflicting signals.
In Fig. 7, we show the difference of using large-scale CSI and full CSI. In this simulation, we set the rate requirement of the interfered user to $3$ bits/s/Hz and assume that the varying frequency of the small-scale fading is one hundred times that of the large-scale fading. The small-scale fading is randomly generated following the Gaussian distribution, i.e., $\mathcal{CN}(0,1)$. Fig. 7 presents one snapshot of the large-scale CSI, which corresponds to $100$ snapshots of the small-scale CSI. From the top subfigure, we can see that, the large-scale-CSI-based scheme cannot guarantee the real-time rate as effectively as the full-CSI-based scheme. This is the compromise of using the large-scale CSI. In return, from the bottom subfigure we can see that the large-scale-CSI-based scheme makes adjustments in large-scale intervals, while the full-CSI-based scheme adjusts the systems in small-scale intervals. This demonstrates that our large-scale-CSI-based scheme significantly lowers the negotiation and adjustment frequency of C&S systems, allowing C&S systems to operate in a loosely cooperated manner.
Algorithm 2 C&S Power Allocation Algorithm According to [26] and Channel-Priority-Based Scheme.
0:  $P_{cmax}$, $P_{rmax}$, $P_{csum}$, $P_{rsum}$, $R_{req}$, $M_{c}$, $M_{r}$, $N_{c}$, and step parameter $\Delta_{P}$.
1:  $\mathbf{P}_{c}=\min\{\frac{P_{csum}}{M_{c}},P_{cmax}\}\mathbf{I}_{M_{c}}$, $\mathbf{P}_{r}=\min\{\frac{P_{rsum}}{M_{r}},P_{rmax}\}\mathbf{I}_{M_{r}}$.
2:  repeat
3:     Calculate $v^{*}$ and $\bar{R}_{ap}(\mathbf{P}_{c},\mathbf{P}_{r},v^{*})$ according to (15) and (14).
4:     if $\bar{R}_{ap}(\mathbf{P}_{c},\mathbf{P}_{r},v^{*})<R_{req}$ then
5:        The scheme proposed in [26]:$p_{r_{i}}=\max\{p_{r_{i}}-\frac{1}{M_{r}}\Delta_{P},0\},\forall i$ The channel-priority-based scheme:$p_{r_{i}}=\max\{p_{r_{i}}-\frac{l_{r_{i}\rightarrow c}p_{r_{i}}}{\sum\limits_{i=1}^{M_{r}}l_{r_{i}\rightarrow c}p_{r_{i}}}\Delta_{P},0\},\forall i$
6:     else if $\bar{R}_{ap}(\mathbf{P}_{c},\mathbf{P}_{r},v^{*})>R_{req}$ then
7:        The scheme proposed in [26]:$p_{c_{i}}=\max\{p_{c_{i}}-\frac{1}{M_{c}}\Delta_{P},0\},\forall i$ The channel-priority-based scheme:$p_{c_{i}}=\max\{p_{c_{i}}-\frac{\sum\limits_{j=1}^{M_{r}}l_{c_{i}\rightarrow r_{j}}p_{c_{i}}}{\sum\limits_{i=1}^{M_{c}}\sum\limits_{j=1}^{M_{r}}l_{c_{i}\rightarrow r_{j}}p_{c_{i}}}\Delta_{P},0\},\forall i$
8:     end if
9:  until $\bar{R}_{ap}(\mathbf{P}_{c},\mathbf{P}_{r},v^{*})=R_{req}$
9:  $\mathbf{P}_{c}$ and $\mathbf{P}_{r}$.
In Fig. 8, we compare the proposed scheme with the scheme proposed in [26], the channel-priority-based scheme, and the unilateral schemes of optimizing $\mathbf{P}_{c}$ or $\mathbf{P}_{r}$ only. The scheme proposed in [26] considered the power allocation between one BS and one radar. Since we consider a multi-BS and multi-radar network, the equal power allocation is used among BSs and radars as the initial allocation, i.e., $\mathbf{P}_{c}=\min\{\frac{P_{csum}}{M_{c}},P_{cmax}\}\mathbf{I}_{M_{c}}$ and $\mathbf{P}_{r}=\min\{\frac{P_{rsum}}{M_{r}},P_{rmax}\}\mathbf{I}_{M_{r}}$, and the step parameter is set as $\Delta_{p}=0.5$ w. Then, per the conclusion drawn in [26], that the optimal power allocation must be that one of C&S systems uses up its available power and the other side reserves power, we thus test the ergodic-rate condition to decide which side should reserve its power. As for the channel-priority-based scheme, we also use equal power allocation as its initial allocation and leverage large-scale CSI of the interference channels to adjust C&S power. Their detailed algorithms are summarized in Algorithm 2. From Fig. 8, we can see that the proposed scheme achieves the highest detection probability. This is because the proposed scheme leverages the large-scale CSI and intelligently allocates more power to the BSs and radars with favorable serving channels and unfavorable interference channels. This amplifies the positive effects of power for their own systems while constraining the negative effects on the opposite system. In contrast, the scheme proposed in [26] is blind to the large-scale CSI, treating all BSs and radars equally in the power allocation. This uniform allocation cannot avoid signal collisions, especially when the transmitting equipment on one side and the receiving equipment on the other side are in close proximity. Only in the special case of one BS and one radar, the scheme proposed in [26] gives the optimal solution. As for the channel-priority-based scheme, when $R_{req}\leqslant 5$ bits/s/Hz, its performance is close to that of the proposed scheme. However, when $R_{req}>5$ bits/s/Hz, the performance of the channel-priority-based scheme experiences severe degradation. Under the initial equal power allocation, the ergodic rate of the interfered user is $5.32$ bits/s/Hz. Therefore, when $R_{req}\leqslant 5$ bits/s/Hz, the BS power is reduced. Under the channel-priority-based scheme, the BS causing stronger interference to radars has its power reduced more. In our topology, the BS causing stronger interference to radars has a poorer channel to the user. Reducing its power greatly lowers the interference to radars while causing small impacts on the user. In contrast, when $R_{req}>5$ bits/s/Hz, the radar power is reduced. Under the channel-priority-based scheme, the radar causing stronger interference to the user has its power reduced more. In our topology, the radar that has the best interference channel to the user is the radar that suffers from the highest communication interference. Although reducing its power greatly reduces the interference, the detection performance of this radar is severely impacted, and consequently, impairs the overall detection performance. Therefore, the power allocation between C&S systems cannot simply consider one factor in isolation. Our proposed scheme provides a solution to consider two-sided impacts of the power in the spectrum-sharing context.
@@
As for the unilateral schemes of optimizing $\mathbf{P}_{c}$ or $\mathbf{P}_{r}$ only. The equal power allocation was used for the not optimized side, i.e., $\mathbf{P}_{r}=\min\{\frac{P_{rsum}}{M_{r}},P_{rmax}\}\mathbf{I}_{M_{r}}$ for optimizing $\mathbf{P}_{c}$ and $\mathbf{P}_{c}=\min\{\frac{P_{csum}}{M_{c}},P_{cmax}\}\mathbf{I}_{M_{c}}$ for optimizing $\mathbf{P}_{r}$. The results of the optimized side are obtained by Algorithm 1. From Fig. 8, we can see that, the efficacy of unilateral schemes is constrained by limited DoFs. Only optimizing $\mathbf{P}_{c}$ fails to provide solutions when $R_{req}\geqslant 6$ bits/s/Hz. We provide a numerical explanation for this phenomenon. When the radar side does not control its transmit power, there is around a tenfold increase in the interference-plus-noise suffered by the user, i.e., from $-107$ dBm to $-98.9$ dBm. Under this condition, the limit ergodic rate is $5.66$ bits/s/Hz, which is obtained by solving the ergodic-rate maximization problem [53]. This indicates that when the radar side does not control its power, no matter how the power is allocated among BSs, the ergodic rate of the interfered user can not exceed $5.66$ bits/s/Hz. This explains why the curve of optimizing $\mathbf{P}_{c}$ only is missing in the region of $R_{req}\geqslant 6$ bits/s/Hz. Therefore, the joint power allocation is more effective in providing scalable C&S services.
@@
In Fig. 9, we show the impact of the power budgets, i.e., $P_{csum}$ and $P_{rsum}$, on C&S performance. It is evident that the detection performance improves with more power resources. In addition, we find that the curves under different $P_{csum}$ overlap when the ergodic rate requirement of the interfered user, i.e., $R_{req}$, is low, and the curves under different $P_{rsum}$ overlap when $R_{req}$ is high. The reason is that, when $R_{req}$ is low, the communication side can meet the ergodic rate requirement without utilizing its full power. To maximize the radar SINR, the communication side reserves its power and the sensing side operates at full capacity. Consequently, the curves under different $P_{csum}$ overlap. In contrast, when $R_{req}$ is high, the sensing side has to control its power and the communication side works at full capacity. This leads to the curves of different $P_{rsum}$ overlapping in the high $R_{req}$ region. We further find that, in spectrum sharing scenarios where two systems are mutually restricted, the optimal solution lies at the boundary of the feasible set but not within it. In other words, one system must exhaust its resources in the optimal solution.
In Fig. 10, we show the impact of radar antenna directivity, depicted by the half-power beam width, i.e., $\theta_{\text{3dB}}$, on the system performance. It can be seen that the detection probability experiences severe degradation as the half-power beam width increases. This is because the larger beam width leads to more power leakage to the interfered user and also opens a larger hole to receive BS interference. In addition, the results show that the gap between the proposed scheme and the scheme proposed in [26] is more obvious when $\theta_{\text{3dB}}$ is large. The reason is that, when $\theta_{\text{3dB}}$ is small, the C&S signals are nearly decoupled by directional beams. The improved space left for the proposed scheme is restricted compared to [26]. In contrast, when $\theta_{\text{3dB}}$ is large, C&S signals are mutually coupled. The proposed scheme functions like directive beams to separate C&S signals in the spatial domain. Consequently, it significantly outperforms the scheme proposed in [26]. In fact, the proposed large-scale-CSI-based scheme works to optimize the spatial distribution of power to mitigate C&S interference.
Figure 11: Zero-Doppler shift and max-Doppler shift cuts of the auto- and cross- ambiguity functions.
Finally, we discuss the impacts of the delay and Doppler shift errors on the radar detection. Specifically, we apply the single-band chirp waveforms as the radar waveforms: $u_{1}(t)=\frac{1}{\sqrt{\tau}}e^{j(\pi B\frac{t^{2}}{\tau}-\pi Bt)},u_{2}(t)=\frac{1}{\sqrt{\tau}}e^{j(-\pi B\frac{t^{2}}{\tau}+\pi Bt)},t\in[0,\tau]$ [51]. The auto- and cross- ambiguity functions of the above two waveforms are given by
$\chi_{1,1}(\Delta{\tau},\Delta{f_{d}})=$
[…]
where $g(\Delta{f_{d}})\triangleq\frac{B}{\tau}+\Delta{f_{d}}$ and $B$ is the waveform bandwidth.
$\chi_{1,2}(\Delta\tau,\Delta{f_{d}})=$
[…]
where $g(\Delta\tau)\triangleq e^{j(\pi B\frac{\Delta\tau^{2}}{\tau}+\pi B\Delta\tau)}.$ We set the pulse duration to $\tau=2$ us, the bandwidth to $B=5$ MHz, and the Doppler shift to $f_{d}=2000$ Hz, corresponding to an approximate velocity of $v\approx 214$ m/s. In Fig. 11, we present the zero-Doppler shift and max-Doppler shift cuts of the auto- and cross- ambiguity functions. It can be seen that the curves representing the zero-Doppler shift and max-Doppler shift, i.e., $\Delta f_{d}=0$ Hz and $\Delta f_{d}=2000$ Hz, are nearly overlapped. In contrast, the ambiguity functions change rapidly with delay errors. This indicates that the timing inaccuracy is the critical factor that impacts the matched filtering. To evaluate this impact, we present the detection performance of the proposed scheme under different delay and Doppler shift errors in Fig. 12. It can be seen that the curves representing different Doppler shift errors are closely overlapped while the curves associated with different delay errors demonstrate a noticeable discrepancy, with a 0.5% delay error bringing 6% detection accuracy degradation. Therefore, aligning delay discrepancies emerges as a pivotal consideration for the distributed radar sensing. Potential solutions include incorporating delay-insensitive detectors and advanced delay compensation algorithms within our framework, which deserve our further studies.
Figure 12: The radar detection performance under different delay and Doppler shift errors, varies with the ergodic rate requirement.